OpenStudy (anonymous):

Use the FOIL method to find the product of the binomials. (x + 7)(3x - 2) A. 3x2 + 21x - 14 B. 3x2 + 19x - 14 C. 3x2 - 2x - 14 D. 3x2 - 23x - 14

OpenStudy (anonymous):

\[(x + 7)(3x - 2) = (x \times 3x) + (7\times-2) + (x \times-2) + (7\times3x) = 3x ^{2} -14 -2x + 21x = 3x ^{2} + 19x -14\] i.e. option B.